OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and parallel to y = 2x + 4. y = -2x + 7 y = 2x - 3 y = 2x - 7 y = -2x + 3

OpenStudy (anonymous):

What is the slope of the first line?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

well the equation is in y=mx+b form aka slope intercept form. The m in the equation is the slope. the number in front of x.

OpenStudy (anonymous):

well the equation is in y=mx+b form aka slope intercept form. The m in the equation is the slope. the number in front of x.

OpenStudy (anonymous):

...this is the whole question n answers idk wut da slop is

OpenStudy (anonymous):

the number in front of x is the slope

OpenStudy (anonymous):

so 2?

OpenStudy (anonymous):

yes. so lines that are parallel have the same slope. the new line slope will be 2 also.

OpenStudy (anonymous):

yes. so lines that are parallel have the same slope. the new line slope will be 2 also.

OpenStudy (anonymous):

to find the new line you need the point and the slope. which you now have.

OpenStudy (anonymous):

to find the new line you need the point and the slope. which you now have.

OpenStudy (anonymous):

Now you will use the formula y-y1=m(x-x1)

OpenStudy (anonymous):

so the aswer is the third one??

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):